OpenStudy (anonymous):

MRS RANDOLPH bought a used car in 2004. In 2007 the car was worth 9500 and in 2010 the car was valued at $4000. express the Value,V, of Mrs. Randolph's car in terms of, t, the number of years she has owned it. a)compute the slope of the graph b) write an equation for v in terms of t

OpenStudy (whpalmer4):

You've got two points, (2007, 9500) and (2010, 4000). Find the slope with \[m = \frac{y_2-y_1}{x_2-x_1}\] Then use the formula for a line given slope and a known point: \[y - y_0 = m(x-x_0)\] Use t and v instead of x and y

OpenStudy (kropot72):



OpenStudy (kropot72):

The two 'given' points on the graph are (3, 9500) and (6, 4000) The slope is given by \[slope=\frac{V _{1}-V _{2}}{t _{1}-t _{2}}\]

OpenStudy (whpalmer4):

Yeah, sorry, I didn't convert to years of ownership.